### v2

# DeepLag: Discovering Deep Lagrangian Dynamics for Intuitive Fluid Prediction

Abstract
Accurately predicting the future fluid is vital to extensive areas such as meteorology, oceanology, and aerodynamics. However, since the fluid is usually observed from an Eulerian perspective, its moving and intricate dynamics are seriously obscured and confounded in static grids, bringing thorny challenges to the prediction. This paper introduces a new Lagrangian-Eulerian combined paradigm to tackle the tanglesome fluid dynamics. Instead of solely predicting the future based on Eulerian observations, we propose DeepLag to discover hidden Lagrangian dynamics within the fluid by tracking the movements of adaptively sampled key particles. DeepLag utilizes the proposed where the Lagrangian movement of the tracked particles is inferred from Eulerian observations, and their accumulated Lagrangian dynamics information is incorporated into global Eulerian evolving features to guide future prediction respectively. Tracking key particles not only provides a transparent and interpretable clue for fluid dynamics but also makes our model free from modeling complex correlations among massive grids for better efficiency. Experimentally, DeepLag excels in three challenging fluid prediction tasks covering 2D and 3D, simulated and real-world fluids.
1 Introduction
Fluids, characterized by a molecular structure devoid of resistance to external shear forces, readily undergo deformation even under minimal force, thereby typically presenting extremely complex dynamics [10]. Consequently, the solvability of theorems for fluid mechanics, such as Navier-Stokes equations, is confined to a limited subset of flows due to the complex form and intricate multiphysics interactions [38]. In practical applications, computational fluid dynamics (CFD) is widely used to infer future fluid by numerical simulations. However, it suffers from substantial computation costs due to the curse of dimensionality [40]. Till now, accurately predicting future dynamics is still a challenging problem. Recently, benefiting from the tremendous non-linear modeling capability, deep models [8, 28, 22] have emerged as a promising tool for fluid prediction, which can be trained with CFD simulations or real-world observations and perform as a fast surrogate model during inference.
A booming direction for deep fluid prediction is learning deep models to solve governing partial differential equations (PDEs) [45]. However, most of these methods [48, 29, 22, 21] attempt to capture fluid dynamics from the Eulerian perspective, which means modeling spatiotemporal correlations among massive grids remains unchanged over time. From this perspective, the complicated moving dynamics in fluids could be seriously obscured and confounded in static grids, bringing challenges in both computational efficiency and learning difficulties in accurately predicting future fluids.
Apart from the Eulerian method, we notice another approach for elucidating fluid dynamics, the Lagrangian method [13], also known as the tracking method. This method primarily focuses on tracing individual fluid particles by conducting the temporal evolution of the position and velocity of each particle. Unlike the Eulerian methods, the Lagrangian approach describes the fluid dynamics through the moving trajectory of individual fluid particles, offering a more natural and neat representation of fluid dynamics with inherent advantages in capturing intricate flow dynamics. Moreover, it helps get around the Courant–Friedrichs–Lewy condition [6] for advection, allowing larger timesteps while retaining stability. As shown in Figure 1, we can find that fluid dynamics is much more apparent in the Lagrangian trajectory than the density change on static Eulerian grids.
Building on the two perspectives mentioned earlier, we propose DeepLag as an Eulerian-Lagrangian Recurrent Network. Our aim is to integrate Lagrangian tracking into the deep model, thereby enhancing the dynamics modeling in Eulerian fluid prediction. To achieve this, we present the EuLag Block, a powerful tool that accomplishes Lagrangian tracking and Eulerian predicting at various scales. By leveraging the cross-attention mechanism, the EuLag Block assimilates tracked Lagrangian particle dynamics into the Eulerian field, guiding fluid prediction. It also forecasts the trajectory and dynamics of Lagrangian particles with the aid of Eulerian features. This unique Eulerian-Lagrangian design harnesses the dynamics information captured by Lagrangian trajectories and the fluid-structure feature learned in the Eulerian grid. In our experiments, DeepLag consistently outperforms existing models, demonstrating its state-of-the-art performance across three representative datasets, covering 2D and 3D fluid at various scales. Our contributions are as follows:
Going beyond learning fluid dynamics at static grids, we propose DeepLag featuring the Eulerian-Lagrangian Recurrent Network, which is the first deep fluid prediction model that explicitly combines Eulerian and Lagrangian frameworks from fluid dynamics concisely.
Inspired by Lagrangian mechanics, we present EuLag Block, which can accurately track particle movements and interactively utilize the Eulerian feature and dynamic information in fluid prediction, enabling a better dynamic modeling paradigm.
DeepLag achieves consistent state-of-the art on three representative fluid prediction datasets with superior trade-offs for performance and efficiency, exhibiting favorable practicability.
2 Preliminaries
2.1 Eulerian and Lagrangian methods
Eulerian and Lagrangian descriptions are two fundamental perspectives for describing fluid motion. The Eulerian view, commonly used in practical applications [23], observes fluid at fixed points and records physical quantities, such as density, as a function of position and time, . Thus, future fluid can be predicted by integrating velocity along the temporal dimension and interpolating results to observed grid points [49]. In contrast, the Lagrangian perspective focuses on the trajectory of individual particles, tracking a particular particle with initial position with its displacement at time . This approach reveals the intricate evolution of the fluid by following particle trajectories, making it convenient for describing complex phenomena like vortices, turbulence, and interface motions [39]. Two perspectives are constitutionally equivalent, as bridged by the velocity:
Furthermore, the material derivative that describes the change rate of a physical quantity of a fluid parcel can be written as the sum of the two terms reflecting the spatial and temporal influence on [1], which represent the derivatives on Eulerian domain and Lagrangian convective respectively:
This connection inspires us to incorporate Lagrangian descriptions into dynamics learning with Eulerian data, enabling a more straightforward decomposition of complex spatiotemporal dependencies.
While traditional particle-based (or mixed-representation) solvers demonstrate superior accuracy and adaptability in inferring small-scale phenomena and dealing with nonlinear and irregular boundary conditions, they require computing the acceleration of each particle through physical equations, followed by sequential updates of their velocity and position [31]. This pointwise modeling approach often demands a significant number of points to fully characterize the dynamics of the entire field to meet accuracy requirements. Moreover, the irregular arrangement of particles and difficulty in parallelization result in higher computational costs and challenges with particle interpolation and gridding [12]. This renders particle-based solvers suboptimal compared to Eulerian solvers, particularly in high-dimensional spaces and large-scale simulations. However, DeepLag leverages the strengths of both solvers, eschewing equations and instead utilizing Eulerian information to assist particle tracking directly. This greatly alleviates the pressure on Lagrangian representation, requiring significantly fewer representative particles to aggregate the dynamics of the entire field.
2.2 Neural Fluid Prediction
Since traditional CFD methods often require hours or even days for precise simulations [41], deep models have been widely explored as efficient surrogate models that can provide near-instantaneous predictions. These neural fluid prediction models approximate the solutions of governing fluid equations differently and can be categorized into three mainstream paradigms as in Figure 2(a-c).
ODE-based Generative Models
As depicted in Figure 2(a), these methods [4, 26, 55] encode multivariate fields into a single-variable latent state , on which they model an ODE governing the state function , representing the first-order time derivative, through neural networks. However, the absence of physical meaning and guidance for latent states in evolution leads to error accumulation and a generally short forecasting horizon. A detailed comparison between DeepLag and these models is provided in Appendix B.
Neural PDE Solvers
This branch of methods in Figure 2(b) adopts deep models to learn the mapping from coordinates to correct values and formalize PDE constraints along with initial and boundary conditions as the loss function of deep models [48, 29, 46, 47]. Though this paradigm can precisely approximate the PDE solution, they usually require exact formalization for coefficients and conditions, limiting their generality and applicability to real-world fluids that are usually partially observed [35]. Plus, the Eulerian input disables them from handling Lagrangian descriptions.
Neural Operators for PDE
Recently, a new paradigm, illustrated in Figure 2(c), has emerged where deep models learn the neural operators between input and target functions, e.g., past observations to future fluid predictions. Various neural operators have significantly advanced fluid prediction following the DeepONet [22], which directly approximates mappings between equation parameters and solutions. For Eulerian grid data, models based on U-Net [32] and ResNet [15] structures have been proposed [30, 27, 17], as well as variants addressing issues like generalizing to unseen domains [44], irregular mesh [11], and uncertainty quantification [51]. Transformer-based models [43] enhance modeling capabilities and efficiency by exploiting techniques like Galerkin attention [3], incorporating ensemble information from the grid [14], applying low-rank decomposition to the attention mechanism [20], and leveraging spectral methods in the latent space [52]. Additionally, FNO [21] learns mappings in the frequency domain, and MP-PDE [2] utilizes the message-passing mechanism. For Lagrangian fluid particle data, some CNN-based methods [34, 42] model particle interactions through redesigned basic modules, while GNN-based methods [33, 25] update particle positions using the Encode-Process-Decode paradigm. Despite the progress made by these methods, they are limited to one description and do not organically combine both perspectives.
3 DeepLag
Following the convention [21], we formalize the fluid prediction problem as learning future fluid given past observation, as shown in Figure 2(d). Given a bounded open subset of -dimensional Euclidean space , a fluid with observed quantities can be seen as functions of coordinates in Banach space . and represent the fluid function at two consecutive time steps for the coordinate . Provided step observations , the fluid prediction process can be written as the following autoregressive paradigm:
where and represents the learned mapping. if and if .
Inspired by Eq. (2), we present DeepLag as an Eulerian-Lagrangian Recurrent Network, which utilizes EuLag Block to learn Eulerian features and important Lagrangian dynamics interactively at various scales to address the complex spatiotemporal correlations in fluid prediction. Specifically, we capture the temporal evolving features at fixed points from the Eulerian perspective and the spatial dynamic information of essential particles from the Lagrangian perspective through their movement. By integrating Lagrangian pivotal dynamic information into Eulerian features, we fully model the spatiotemporal evolution of the fluid field over time and motion. Moreover, DeepLag can obtain critical trajectories within fluid dynamics with high computational efficiency by incorporating high-dimensional Eulerian space into lower-dimensional Lagrangian space.
3.1 Overall Framework
It is widely acknowledged that fluids exhibit different motion characteristics across varying scales [3, 52]. In order to capture the intrinsic dynamics information at different scales, we track the trajectories of the key particles on these scales separately and propose an Eulerian-Lagrangian Recurrent Network to realize the interaction between Eulerian and Lagrangian information.
Initializing Lagrangian particles
To better capture the complex dynamics within the fluid field, we sample crucial points from the given observations in the Eulerian perspective to determine the initial position for Lagrangian tracking in the dynamic sampling module. Explicitly, for the first predicting step, i.e. , given observations on all points with channels at the -th scale, we extract their spatial dynamics by a convolutional structure. We then calculate the probability matrix using softmax along spatial dimension:
where consists of a convolutional, activation, and fully-connected layer. We then sampled tracking particles using the probability matrix at the -th scale,
where represents the set of sampled particles.
Eulerian Fluid Prediction with Lagrangian Dynamics
For the -th step, we adopt a learnable embedding layer with multiscale architecture to encode past observations or predictions to obtain the Eulerian representations , where , and denotes the corresponding downsampled observation domain.
At each scale, we integrate EuLag Block to fuse the Lagrangian dynamics, thus directing the evolution of Eulerian features. This interaction also enables Eulerian features to guide the progression of Lagrangian dynamics. For the -th scale, we track key particles over time, with representing their positions and denoting their learned particle dynamics. As shown in Figure 2(d), the positions and dynamics of particles at the -th scale are learned autoregressively using the EuLag Block, which can be written as:
where the scale index is omitted. The EuLag Block optimally leverages the complementary strengths of Eulerian and Lagrangian representations, facilitating enhanced training and mutual refinement between these two perspectives. More details about the implementation of the EuLag Block are expanded in the following subsection.
After getting the output Eulerian features with Lagrangian particle position and dynamics , we further aggregated with at coarser scale by upsampling to be . Eventually, the prediction at step is decoded from with a projection layer. We unfold the implementation of the overall architecture in Appendix A.2.
3.2 EuLag Block
As formalized in Eq. (6), we adopt a recurrent framework to interactively exploit the information from Eulerian and Lagrangian perspectives, which consist of two main components: Lagrangian-guided feature evolving and Eulerian-conditioned particle tracking. For clarity, we omit the scale index in this section. The following quantities are all in the -th scale.
Lagrangian-guided feature evolving
Classical theories [9] and numerical algorithms [31] showed that fluid predictions can be solved by identifying the origin of the fluid parcel and interpolating the dependent variable from nearby grid points. However, without specifying certain PDEs, we cannot explicitly determine the former position of the particle on each Eulerian observed point. Thus, we first adaptively synthesize the Lagrangian dynamics of the tracked particles to guide the evolution of Eulerian features using a cross-attention mechanism. Expressly, we adopt a Lagrangian-to-Eulerian attention, where the Eulerian field serves as queries, and Lagrangian dynamics with particle position is used as keys and values:
where , and stand for linear projection layers. We summarize the above Lagrangian dynamics guided Eulerian Evolution process as .
Eulerian-conditioned particle tracking
Traditional Lagrangian methods rely on interactions among vast quantities of particles to estimate the future fluid fields. However, the high computational cost hinders the application in rapid, deep models. In other data-driven approaches, the sparse sampling of tracking particles is computational-friendly but cannot directly derive Lagrangian dynamics for every particle. Considering the equivalence of the Eulerian and Lagrangian representations indicated in Eq. (1), we propose to learn particle movements based on the Eulerian conditions. Concretely, we utilize another Eulerian-to-Lagrangian cross-attention, where the evolved Eulerian features of the whole fluid field are used to navigate the Lagrangian dynamics of sparse particles:
where represents standard transformer block with as linear projection matrix. The forecasted global Lagrangian dynamic and movement of tracking particles are produced by two transformers with the same query and key projection but different value projections. To better model the dynamic evolution of the particles, we gather local Lagrangian dynamic by employing bilinear interpolation to output Eulerian features on particles , then use a linear function to interact with global dynamics information , naming . Additionally, as the input flow field may have an open boundary, the particles could move out of the observation domain. We check the position of the particles and resample them in case they exit. Similar to the previous Eulerian Evolution process, the above Lagrangian Tracking process is summarized as .
Overall, the EuLag Block can fully utilize the complementary advantages of Eulerian and Lagrangian perspectives in describing fluid dynamics, thereby being better suited to fluid prediction. For more implementation details of the EuLag Block, please refer to Appendix A.3.
4 Experiments
We evaluated DeepLag on three challenging benchmarks, including simulated and real-world scenarios, covering both 2D and 3D, as well as single and multi-physics fluids. Following the previous convention [21], we train DeepLag and the baselines for each task to predict ten future timesteps in an autoregressive fashion given ten past observations. Detailed benchmark information is listed in Table 1. We provide an elaborate analysis of the efficiency, parameter count, and performance difference in section 4.4 and Appendix E. Additionally, more detailed visualizations, besides in later this section, are provided in Appendix H. Furthermore, the trained models are engaged to perform 30 frames extrapolation to examine their long-term stability.
| Datasets | Type | #Var | #Dim | #Space |
| Bounded N-S | Simulation | 1 | 2D | |
| Ocean Current | Real World | 5 | 2D | |
| 3D Smoke | Simulation | 4 | 3D |
Baselines
To demonstrate the effectiveness of our model, we compare DeepLag with seven baselines on all benchmarks, including the classical multiscale model U-Net [32] and advanced neural operators for Navier-Stokes equations: FNO [21], Galerkin Transformer [3], Vortex for 2D image [7], GNOT [14], LSM [52] and FactFormer [20]. U-Net has been widely used in fluid modeling, which can model the multiscale property precisely. LSM [52] and FactFormer [20] are previous state-of-the-art neural operators.
Metrics
Implementations
Aligned with convention and the baselines, DeepLag is trained with relative L2 as the loss function on all benchmarks. We use the Adam [18] optimizer with an initial learning rate of and StepLR learning rate scheduler. The batch size is set to 5, and the training process is stopped after 100 epochs. All experiments are implemented in PyTorch [24] and conducted on a single NVIDIA A100 GPU. Training curves are shown in Appendix D.
4.1 Bounded Navier-Stokes
Setups
In real-world applications, handling complex boundary conditions in predicting fluid dynamics is indispensable. Thus, we experiment with the newly generated Bounded Navier-Stokes, which simulates a scenario where some colored dye flows from left to right through a 2D pipe with several fixed pillars as obstacles inside. Details about this benchmark can be found in Appendix C.1.
| Model | Relative L2 () | |
| 10 Frames | 30 Frames | |
| U-Net [32] | 0.0618 | 0.1038 |
| FNO [21] | 0.1041 | 0.1282 |
| Galerkin Transformer [3] | 0.1084 | 0.1369 |
| Vortex [7] | 0.1999 | NaN |
| GNOT [14] | 0.1388 | 0.1793 |
| LSM [52] | 0.0643 | 0.1020 |
| FactFormer [20] | 0.0733 | 0.1195 |
| DeepLag (Ours) | 0.0543 | 0.0993 |
| Promotion | 13.8% | 2.7% |
Quantitive results
As shown in Table 2, DeepLag achieves the best performance on Bounded Navier-Stokes, demonstrating its advanced ability to handle complex boundary conditions. In comparison to the previous best model, DeepLag achieves a significant 13.8% (Relative L2: 0.0618 v.s. 0.0544) and 2.7% (Relative L2: 0.1020 v.s. 0.0993) relative promotion on short and long rollout. The timewise error curves of all the models are also included in Figure 4. We can find that DeepLag presents slower error growth and excels in long-term forecasting. This result may stem from the Lagrangian-guided fluid prediction, which can accurately capture the dynamics information over time, further verifying the effectiveness of our design.
Showcases
To intuitively present the forecasting skills of different models, we also provide showcase comparisons in Figure 4. We can find that DeepLag can precisely illustrate the vortex in the center of the figure and the Kármán vortex phenomenon formed behind the upper left pillar. As for U-Net and LSM, although they successfully predicted the position of the center vortex, the error map shows that they failed to predict the density field accurately. In addition, FactFormer deteriorates on this benchmark. This may be because it is based on spatial factorization, which is unsuitable for irregularly placed boundary conditions. These results further highlight the benefits of Eulerian-Lagrangian co-design, which can simultaneously help with dynamic and density prediction.
4.2 Ocean Current
| Model | Relative L2 () | |
| 10 Days | 30 Days | |
| U-Net [32] | 0.0185 | 0.0297 |
| FNO [21] | 0.0246 | 0.0420 |
| Galerkin Transformer [3] | 0.0323 | 0.0515 |
| Vortex [7] | 0.9548 | NaN |
| GNOT [14] | 0.0206 | 0.0336 |
| LSM [52] | 0.0182 | 0.0290 |
| FactFormer [20] | 0.0183 | 0.0296 |
| DeepLag (Ours) | 0.0168 | 0.0257 |
| Promotion | 8.3% | 12.8% |
Setups
Predicting large-scale ocean currents, especially in regions near tectonic plate boundaries prone to disasters such as tsunamis due to intense terrestrial activities, plays a crucial role in various domains. Hence, we also explore this valuable scenario in our experiments. More details about this benchmark can be found in Appendix C.2.
Quantitive results
We report relative L2 for the Ocean Current dataset in Table 3, where DeepLag still achieves the best with 8.3% relative promotion w.r.t. the second-best model. Even in 30 days of forecasting, the number is 12.8%. These results show that DeepLag performs well in real-world, large-scale fluids, which usually involve more inherent stochasticity than simulated data. Moreover, we provide the ACC metric and timewise curve in Appendix F.
Showcases
We presented different model predictions in Figure 5. In comparison to other models, DeepLag exhibits the most minor prediction error. It accurately predicts the location of the high-temperature region to the south area and provides a clear depiction of the Kuroshio pattern [37] bounded by the red box.
Learned trajectory visualization
To reveal the effect of learning Lagrangian trajectories, we visualize tracked particles in Figure 5. We observe that all the particles move from west to east, consistent with the Pacific circulation. Additionally, the tracked particles show distinct moving patterns, confirming their ability to represent complex dynamics. The movement of upper particles matches the sinuous trajectory of the Kuroshio current, demonstrating the capability of DeepLag to provide interpretable evidence for prediction results. Visualizing the tracking points in Lagrangian space instills confidence in the reliability and interpretability of the predictions made by our system.
4.3 3D Smoke
Setups
3D fluid prediction has been a long-standing challenge due to the tanglesome dynamics involved. Therefore, we created this benchmark to describe a scenario in which smoke flows under the influence of buoyancy in a three-dimensional bounding box. For more details, please refer to Appendix C.3.
Quantitive results
Table 4 shows that DeepLag still performs best in 3D fluid. Note that in this benchmark, the canonical deep model U-Net degenerates seriously, indicating that a pure Eulerian multiscale framework is insufficient to model complex dynamics. We also noticed that the Transformer-based neural operators, such as GNOT and Galerkin Transformer, failed in this task. This may be because both of them are based on the linear attention mechanism [19, 54], which may depreciate under massive tokens. These comparisons further highlight the capability of DeepLag to handle high-dimensional fluid.
Showcases
In Figure 6, we compared prediction results on the 3D Smoke dataset. DeepLag performs well in capturing the convection and diffusion of smoke within the bounding box. In comparison, the predictions of U-Net tend to predict the averaged value across various surfaces, resulting in blurred details, which also indicates its deficiency in dynamics modeling. As for LSM and FactFormer, we can find that they present more errors in the smoke boundaries that usually involve complex waves. In contrast, our model excels in overall and cross-section error and performs excellently in predicting subtle flows.
4.4 Model Analysis
Ablations
To verify the effectiveness of detailed designs in DeepLag, we conduct exhaustive ablations in Table 5. In our original design, we track 512 particles (around 3% of Eulerian grids).
| Design | Relative L2 () | Decrease |
| DeepLag () | 0.0543 | - |
| 0.0559 | 2.9% | |
| 0.0553 | 1.8% | |
| 0.0547 | 0.7% | |
| w/o | 0.0556 | 2.4% |
| w/o | 0.0547 | 0.7% |
| w/o Learnable Sampling | 0.0552 | 1.6% |
The experiments indicate that further increasing the number of particles yields marginal performance improvements. Therefore, we opt for to strike a balance between efficiency and performance. In addition, we can conclude that all components proposed in this paper are indispensable. Especially, the lack of interaction between the Eulerian and Lagrangian space will cause a severe drop in accuracy, highlighting the dual cross-attention that exploits Lagrangian dynamics has a positive influence on the evolution of Eulerian features. Besides, rather than uniformly sampling particles, sampling from a learnable probability matrix also provides an upgrade (refer to Appendix G for visual results). The above results provide solid support to our motivation in tracking essential particles and utilizing Eulerian-Lagrangian recurrence, further confirming the merits of our model.
Efficiency analysis
We also include the efficiency comparison in Figure 7. DeepLag hits a favorable balance between efficiency and performance by simultaneously considering performance, model parameters, and running time, demonstrating superior performance with significantly less memory than GNOT and LSM, thereby minimizing storage complexity. Standard U-Net has a large number of parameters, and Transformers have quadratic memory, so in large-scale or complex fluid prediction scenarios, using linear complexity attention mechanisms like ours is necessary. This explains why, although U-Net and LSM are good at Bounded Navier-Stokes, they will worsen when it comes to complex fluid dynamics, such as the Ocean Current and the 3D Smoke benchmark.
5 Conclusions and Limitations
To tackle intricate fluid dynamics, this paper presents DeepLag by introducing the Lagrangian dynamics into Eulerian fluid, which can provide clear and neat dynamics information for prediction. A EuLag Block is presented in the Eulerian-Lagrangian Recurrent Network to utilize the complementary advantages of Eulerian and Lagrangian perspectives, which brings better particle tracking and Eulerian fluid prediction. DeepLag excels in complex fluid prediction with an average improvement of nearly 10% across three carefully selected complex benchmarks, even in 3D fluid, and can also provide interpretable evidence by plotting learned Lagrangian trajectories. However, the number of tracking particles in DeepLag is appointed as a fixed hyperparameter that needs to be adjusted for specific scenarios, leaving space for future exploration.
Appendix A Implementation Details
This section provides the implementation details of DeepLag, including the configurations of model hyperparameters and the concrete design of modules.
A.1 Hyperparameters
Detailed model configurations of DeepLag are listed in Table 6. Zero-padding is only used in the Ocean Current dataset to ensure the exact division in downsampling.
| Model Designs | Hyperparameters | Values |
| Number of observation steps | ||
| Number of scales | ||
| Eulerian-Lagrangian | Sample Points at each scale | |
| Recurrent Network | Downsample Ratio | |
| Channels of each scale | ||
| Paddings for Ocean Current dataset | ||
| EuLag Block | Heads in Cross-Attention | 8 |
| Channels per head in Cross-Attention | 64 |
A.2 Sampling and multiscale architecture in overall framework
A.2.1 Sampling to initialize Lagrangian particles
At the first predicting step, the position of Lagrangian particles to track is initialized by the dynamic sampling module consisting of dynamics extracting and sampling.
The to extract dynamics
Given the Eulerian input function at the -th scale, the operation in Eq. (4) is to extract the local dynamics around each Eulerian observation point with , and layers, which can be formalized as follows:
Here, the output channel of the outermost is 1.
The to select key particles
Given the probability distribution matrix and the number of particles to sample at the -th scale, we choose the particles for further tracking by without replacement:
A.2.2 Multiscale architecture
Multiscale modeling is utilized in DeepLag as represented in Figure 2(d), where we need to maintain multiscale deep Eulerian features as follows:
Encoder
Given Eulerian fluid function at the -th time step, at previous time steps and the 0-1 boundary-geometry mask where 1 indicates the border and the unreachable area (like pillars in Bounded Navier-Stokes), the operation is to project original fluid properties in physical domain to deep representations with linear layer and position embedding, which can be formalized as follows:
Decoder
Given evolved Eulerian deep representations in the finest scale at the -th time step, the operation is to project deep representations back to predicted fluid properties with two linear layers and a activation [16], which can be formalized as follows:
Downsample
Given Eulerian deep representations at the -th scale, the operation is to concentrate local information of deep representations into a smaller feature map at the -th scale with and layers, which can be formalized as follows:
Upsample
Given the evolved Eulerian deep representations at the -th scale and at the -th scale, respectively, the operation is to fuse information on corresponding position between two adjacent scales of deep representations into a feature map at the -th scale with operation and layers, which can be formalized as follows:
A.3 EuLag Block
The process
As we stated in subsection 3.2, the Lagrangian-guided Eulerian feature evolving process, short as , aggregates information from Lagrangian description to guide the update of Eulerian field with a single Transformer layer at each scale:
where is described as Eq. (7). Moreover, we use the pre-normalization [53] technique for numerical stability.
The process
Similar to above, we acquire the new particle position and the global dynamics in the Eulerian-conditioned particle tracking by a Eulerian-Lagrangian cross-attention:
where . Then, we interpolate from by to be the local dynamics and Aggrerate it with the global dynamics using a Linear layer:
A.4 Metrics
Relative L2
We use the relative L2 as the primary metric for all three tasks. Compared to MSE, Relative L2 is less influenced by outliers and is more robust. For given steps 2D predictions or 3D predictions and their corresponding ground truth of the same size, the relative L2 can be expressed as:
where represents the L2 norm.
Appendix B Comparison Between DeepLag and Neural ODE-based Methods
DeepLag differs significantly from ODE-based approaches like Neural ODE [4], as highlighted in the comparison Table 7:
| Feature | DeepLag | Neural ODE |
| ODE Specification | Not required | Explicitly specified |
| Computational Paradigm | Data-driven | Model-driven |
| Attention Mechanism | Utilizes attention | Typically not utilized |
| Integration of Lagrangian Dynamics | Integrated | Operates solely in Euler space |
| Input-Output Mapping | Complex, high-dimensional PDE | Simple, single-variable ODE |
While Neural ODE methods explicitly specify ODEs, fluid dynamics is governed by multi-variable PDEs, rendering the ODE formulation inadequate. Moreover, DeepLag is data-driven and does not require explicit ODE specification. Notably, we are the first to employ attention mechanisms for computing Lagrangian dynamics, which closely resemble operators in both deep learning and numerical Lagrangian methods. Additionally, DeepLag integrates both Eulerian and Lagrangian frameworks, whereas ODE-based methods operate solely in the Eulerian space. Finally, DeepLag models the complex mapping of high-dimensional spatiotemporal PDEs, which is significantly more intricate than the simpler processes modeled by single-variable (either temporal or spatial) ODEs.
Appendix C More Details about the Benchmarks
C.1 Bounded Navier-Stokes
Here we provide some details about the benchmark Bounded Navier-Stokes. Just as its name implies, the motion of dye is simulated from the Navier-Stokes equation of incompressible fluid. 2000 sequences with spatial resolution of are generated for training and 200 new sequences are used for the test. We supplement important details indicating the difficulty of the Bounded Navier-Stokes dataset as follows:
About the Reynolds number
The Reynolds number of the dataset is 256. At this Reynolds number, attached vortices dissipate and form a boundary layer separation. The downstream flow field behind the cylinder becomes unsteady, with vortices shedding periodically on both sides of the cylinder’s rear edge, resulting in the well-known phenomenon of Kármán vortex street [50]. Additionally, due to the presence of multiple cylinders within the flow field and obstruction from downstream cylinders, more complex flow phenomena than flow around a cylinder occur, challenging the model’s capacity.
Differences of data sequences
Our data generation method involves running simulations for over steps after setting the initial conditions of the flow field. We then randomly select a starting time step and extract several frames as an example, which are further randomly divided into training and testing sets. Due to the highly unsteady nature of the flow field, the flow patterns observed by the model appear significantly different.
Numerical method used in data generation
We utilized the Finite Difference Method (MAC Method) with CIP (Constrained Interpolation Profile) as the Advection Scheme for numerical simulations, implemented by [36]. This high-order interpolation-constrained format effectively reduces numerical dissipation, enhancing the accuracy and reliability of numerical simulations.
C.2 Ocean Current
Some important details related to the Ocean Current benchmark are attached here. Learning ocean current patterns from data and providing long-term forecasts are of great significance for disaster prevention and mitigation, which motivates us to focus on this problem.
The procedures to make this dataset are as follows. First, we downloaded daily sea reanalysis data [5] from 2011 to 2020 provided by the ECMWF and selected five basic variables on the sea surface to construct the dataset, including velocity, salinity, potential temperature, and height above the geoid, which are necessary to identify the ocean state. Then, we crop a sub-area on the North Pacific from the global record, corresponding to a 375km625km region. In total, this dataset consists of 3,653 frames, where the first 3000 frames are used for training and the last 600 frames are used for testing. The training task is to predict the future current of 10 days based on the past 10 days’ observation, after which we performed 30 days of inference with the trained model to examine the long-term stability of DeepLag.
C.3 3D Smoke
To verify our model effectiveness in this complex setting of the high-dimensional tanglesome molecular interaction, we also generate a 3D fluid dataset for the experiment. This benchmark consists of a scenario where smoke flows under the influence of buoyancy in a three-dimensional bounding box. This process is governed by the incompressible Navier-Stokes equation and the advection equation of fluid. 1000 sequences are generated for training, and 200 new samples are used for testing. Each case is in the resolution of .
Appendix D Training Curves
We provide training curves on Bounded Navier-Stokes, Ocean Current and 3D Smoke datasets in Figure 8. We can observe that DeepLag presents favorable training robustness and converges the fastest on the Bounded Navier-Stokes dataset.
Appendix E Analysis on the Parameter Count and Performance Difference
In Table 1 in our paper, we present three different datasets with variations in data type, number of variables, number of dimensions, and spatial resolution. Specifically, the 2D fluid and 3D fluid represent entirely different dynamic systems, so it is normal for different baselines to perform inconsistently across different benchmarks. In the following results, patch_size of all models are set to 1 for fair comparison. For instance, U-Net performs well on datasets with distinct multiscale attributes (such as the Bounded Navier-Stokes dataset), while transformer-based methods excel on datasets with broad practical ranges (like the Ocean Current dataset) or high dimensions (such as the 3D Smoke dataset), where attention mechanisms can effectively model global features. Hence, our model’s ability to handle multiscale and global modeling simultaneously highlights the challenge of achieving consistent state-of-the-art performance. Below are the parameter statistics for DeepLag and all baseline models across each benchmark, along with some experiments we conducted to ensure parity in the parameter count for each model.
E.1 Analysis on Bounded Navier-Stokes benchmark
The parameter quantities for each model on the Bounded Navier-Stokes benchmark are shown in Table 8. When attempting to further increase the parameter count of smaller models for fair comparison, we encountered CUDA Out-of-memory errors. Specifically, these errors occurred when we increased the Dynamic_net_layer of Vortex from 4 to 50, resulting in a parameter count of 171,761, but encountered CUDA Out of Memory issues.
E.2 Analysis on Ocean Current benchmark
The parameter quantities for each model on the Ocean Current dataset are summarized in Table 9. When attempting to further increase the parameter count of smaller models for fair comparison, we encountered CUDA Out-of-memory errors. Specifically, these errors occurred when we increased the Dynamic_net_layer of Vortex from 4 to 50, resulting in a parameter count of 171,761, but encountered CUDA Out of Memory issues. Similarly, for GNOT, increasing the latent_dim from 64 to 96 resulted in a parameter count of 1,659,179, yet this configuration encountered CUDA Out of Memory errors.
E.3 Analysis on 3D Smoke benchmark
The parameter quantities for each model on the 3D Smoke dataset are shown in Table 10. When experimenting with increasing the parameter count of certain models for fair comparison, we encountered CUDA Out-of-memory errors. Specifically, these errors occurred when we increased the latent_dim of GNOT from 64 to 96, resulting in a parameter count of 1,450,768 for GNOT, which encountered CUDA Out of Memory issues. Similarly, when adjusting the encoder_transformer_layer of FactFormer from 3 to 13, the parameter count reached 7,382,084, yet this configuration encountered CUDA Out of Memory errors.
Figure 7 in our paper, along with the provided tables above, illustrate our rigorous comparison and efforts to standardize model parameters for a fair comparison. However, due to excessive GPU memory consumption of intermediate results in some baseline models, we could not conduct direct performance comparisons under matched parameter conditions. The ability of our model to efficiently utilize GPU memory resources is a valuable aspect of practical applications.
Appendix F ACC Metric on Ocean Current
Latitude-weighted Anomaly Correlation Coefficient
In meteorology, directly calculating the correlation between predictions and ground truth may obtain misleadingly high values because of the seasonal variations. To subtract the climate average from both the forecast and the ground truth, we utilize the Anomaly Correlation Coefficient to verify the forecast and observations. Moreover, since the observation grids are equally spaced in latitude, the size of the different grids is related to the latitude and thus, we calculate the latitude-weighted Anomaly Correlation Coefficient, which can be formalized as:
where represents a certain observed variable, is the prediction of ground truth at position and forecast time . represents the difference between and the climatology , that is, the long-term mean of observations in the dataset. , where and is the latitude of the -th row of output.
ACC result on the Ocean Current benchmark
Notably, DeepLag also excels in the ACC metric as shown in Table 11, which can better quantify the model prediction skill. As shown in the timewise ACC curve, DeepLag consistently achieves the highest ACC and holds more significant advantages in long-term prediction. Since ACC is calculated with long-term climate statistics, further improvements become increasingly challenging as it increases, which highlights the value of DeepLag.
Appendix G Visual Results for learnable sampling
To demonstrate the effectiveness of our learnable probability, we visualize its distribution with respect to the pointwise variance of vorticity, which is directly proportional to the local complexity of fluid dynamics, in Figure 9. It is evident that our sampling module tends to prioritize regions with higher dynamical complexity, such as the wake flow and Karman vortex [50] behind the pillar. The showcase in Figure 1 also demonstrates that the tracked particle can well present the dynamics of a certain area. This observation underscores that our design effectively guides the model to track the most crucial particles, enhancing its performance.
Appendix H More Showcases
Appendix I Result of Long-term Prediction
I.1 Reason for predicting 10 steps
Predicting the last 10 steps from the first 10 steps input is a convention (refer to FNO where is the setting, and later several baselines using the NS dataset also followed this setting). We did this to follow the convention and for ease of comparison. Additionally, the Ocean Current dataset has a one-day interval between every two frames, and predicting 10 days ahead is already a long horizon.
I.2 Results of long-term rollout
We conducted experiments for long-term (extrapolation beyond hundreds of frames) predictions. Accurately, we utilize trained 10-frame prediction models to perform 100-frame prediction. Our reason for not directly training a model in an autoregressive paradigm to predict the next 100 frames is due to insufficient memory capacity and the issue of gradient explosion or vanishing. The results are as follows. The best result is bold and the second is underlined. The reason we did not run the 3D Smoke dataset is that it is very challenging to load 110 frames of large 3D data at once, which overwhelms our machine. Extra video results are in supplementary materials, which effectively illustrate the performance and consistency of our model’s predictions.
I.2.1 Bounded Navier-Stokes
Quantitive results
As depicted in Table 12, the DeepLag model still outperforms the strong baselines on the Bounded Navier-Stokes dataset in predicting 100 frames into the future. With a Relative L2 of 0.1493, DeepLag achieves superior performance compared to all baselines. Additionally, the performance trends of all models are visually illustrated in Figure 16 - 20, revealing DeepLag’s ability to maintain lower error growth rates over time, particularly in long-term predictions. This outcome suggests that the Lagrangian particle-based approach adopted by DeepLag effectively captures dynamic information, contributing to its robust forecasting capability in fluid dynamics modeling.
Showcases
To visually evaluate the predictive capabilities of our models on long-term, we present a showcase of last frame comparisons and time-wise prediction in Figure 16 - 20, illustrating the long-term rollout performance on the Bounded Navier-Stokes dataset. Notably, DeepLag demonstrates remarkable accuracy in capturing complex flow phenomena, accurately predicting the formation and evolution of vortices, particularly the Kármán vortex street behind the upper left pillar. In contrast, U-Net and LSM exhibit moderate success in predicting the central vortex but struggle with accurately reproducing the density distribution of the flow field, as indicated by the error maps. FactFormer, however, shows subpar performance on this benchmark, likely due to its reliance on spatial factorization, which may not effectively handle irregular boundary conditions. These findings underscore the advantages of our Eulerian-Lagrangian co-design approach, which enables simultaneous prediction of dynamics and density, contributing to more accurate and comprehensive fluid modeling and forecasting capabilities. Here are figure showcases of Bounded Navier-Stokes dataset and time-wise prediction with error map:
I.2.2 Ocean Current
Quantitive results
We present a comparison of results for the Ocean Current dataset in Table 13, which includes relative L2, Last frame ACC, and Average ACC metrics. DeepLag maintains its superiority, achieving the lowest relative L2 among all models, with an 8.7% promotion compared to the second-best model. These findings highlight DeepLag’s robust performance in predicting real-world large-scale fluid dynamics, which often exhibit inherent stochasticity. Furthermore, DeepLag outperforms other models in ACC metrics, indicating its superior predictive capability. This is further corroborated by the timewise ACC curve, where DeepLag consistently demonstrates the highest ACC values, particularly in long-term predictions.
| Model | Relative L2 () | Last frame ACC () | Average ACC () |
| U-Net [32] | 0.0764 | 0.0197 | 0.0503 |
| FNO [21] | 0.1675 | 0.0041 | 0.0448 |
| Galerkin-Transformer [3] | NaN | NaN | NaN |
| GNOT [14] | 0.0513 | 0.173 | 0.3586 |
| LSM [52] | 0.0616 | 0.0793 | 0.1650 |
| FactFormer [20] | 0.0460 | 0.2647 | 0.3864 |
| DeepLag (Ours) | 0.0423 | 0.2890 | 0.4041 |
| promotion | 8.7% | 9.2% | 4.6% |
Showcases
To visually assess the long-term forecasting performance of each model, we showcase the last frame predictions along with their errors, and the time-wise prediction errors for each model in Figure 21 - 25. Visually, our DeepLag predictions closely resemble the Ground Truth compared to other models, demonstrating robust long-term extrapolation capabilities and accurate capture of the Kuroshio pattern [37]. It can be observed that FactFormer and LSM exhibit relatively large errors, while GNOT tends to average and loses fine texture details. However, DeepLag does not suffer from these issues.
Here are figure showcases of the Ocean Current dataset:
And time-wise prediction with error map:
I.3 Examination on the turbulent kinetic energy spectrum
In the field of fluid mechanics, simulation results that better adhere to intrinsic physical laws are sometimes more valuable than those with smaller pointwise errors, often reflected in frequency domain analysis. To validate this point and to measure long-term forecasting ability, we introduced a metric on time-averaged turbulent statistics, namely the turbulent kinetic energy spectrum (TKES). Specifically, we computed the MAE and RMSE of TKES for the Bounded Navier-Stokes dataset, which exhibits the most prominent turbulent characteristics, as shown in Table 14, and plotted the line graphs of the error of wave number and TKE, as depicted in Figure 26. It can be observed both numerically and visually that DeepLag consistently outperforms various baselines to different extents.
| Model | MAE () | RMSE () |
| U-Net | 3.888e+06 | 3.349e+07 |
| FNO | 7.364e+06 | 7.191e+07 |
| Galerkin-Transformer | 5.956e+06 | 5.950e+07 |
| GNOT | 1.210e+07 | 1.631e+08 |
| LSM | 3.403e+06 | 3.040e+07 |
| FactFormer | 4.495e+06 | 3.968e+07 |
| DeepLag (Ours) | 3.052e+06 | 2.716e+07 |
| promotion | 11.5% | 11.9% |